OpenStudy (anonymous):

how do you solve the lim x 4 square root of x-2 divided by x-4

geerky42 (geerky42):

You mean this? \(\large \displaystyle\lim_{x\rightarrow4}\dfrac{\sqrt{x-2}}{x-4}\)

OpenStudy (anonymous):

yes

geerky42 (geerky42):

@hartnn

hartnn (hartnn):

but i think the question is \(\large \displaystyle\lim_{x\rightarrow4}\dfrac{\sqrt{x}-2}{x-4}\) is it ?

geerky42 (geerky42):

is -2 in square root or no? @Landva2

hartnn (hartnn):

its outside

hartnn (hartnn):

do you know \(a^2 - b^2 = .. ?\)

hartnn (hartnn):

\(\Large x-4 = (\sqrt x)^2 - 2^2 = ... ?\)

OpenStudy (anonymous):

I am not seeing how that equal 1/4 maybe I am doing something wrong

geerky42 (geerky42):

\(a^2-b^2=(a-b)(a+b)\) So \((\sqrt x)^2-2^2=\ldots?\)

geerky42 (geerky42):

Do you know what to do? @Landva2 we are trying to "factor" denominator